OpenStudy (anonymous):

a vessel is of the shape of a cone.radiusof the broader end is2.1cm and height is 20cm find volume of the vessel?

OpenStudy (anonymous):

what you need is a formula

OpenStudy (anonymous):

i have calculated, wanna recheck my solution Given - r = 2.1 cm and h = 20 cm Volume of cone = 1/3 pie * r *r *h =1/3 * 22/7 *2.1 *2.1 *20 =1940.4/21 = 92.4 cm^3

OpenStudy (anonymous):

or you want to know how to integral ?

OpenStudy (anonymous):

did i solve right, plse tell me

OpenStudy (anonymous):

can anyone tell me vat i did is right or wrong

OpenStudy (anonymous):

yes your formula is right